# Dissecting the Hubble tension: Insights from a diverse set of Sound Horizon-free measurements

### v1

@@ Acknowledgements.
This research was supported by COST Action CA21136 - Addressing observational tensions in cosmology with systematics and fundamental physics (CosmoVerse), supported by COST (European Cooperation in Science and Technology).Appendix A Complete Data Tables
@@
| 1. Distance Ladder | 29 | 72.74 | 0.40 | 0.74 | 67.8–76.9 |
| 2. Local One-Step (CDM) | 33 | 67.18 | 0.90 | 0.85 | 59.1–86.2 |
| 3. Pure Local One-Step | 15 | 70.38 | 1.00 | 0.78 | 65.9–75.5 |
| 4. CMB Sound Horizon Free | 6 | 68.71 | 0.63 | 0.36 | 61–69.5 |
| Combined One-Step (2+3+4) | 54 | 68.67 | 0.46 | 0.85 | 59.1–86.2 |
@@ A.1 Distance Ladder Measurements: Extended Information
Table 9 provides extended information for all 29 distance-ladder-dependent measurements. We include the geometric anchor(s) employed, the primary distance indicator, the secondary distance indicator (if applicable), and the reported value with uncertainty.
@@
| 1 | Huang (2019) [Huang2020] | NGC 4258, MW, LMC | Mira | SNe Ia | ||
| 2 | de Jaeger (2020) [deJaeger2020] | LMC, MW | Cepheids/TRGB | SNe II | ||
| 3 | Kourkchi (2020) [Kourkchi2020] | Cepheids, TRGB | Tully–Fisher | — | ||
| 4 | Khetan (2021) [Khetan2021] | Cepheids, TRGB | SBF | SNe Ia | ||
| 5 | Blakeslee (2021) [Blakeslee2021] | NGC 4258, LMC | SBF | SNe Ia | ||
| 6 | Freedman (2021) [Freedman2021] | LMC (DEB), MW | TRGB | SNe Ia | ||
| 7 | Dhawan (2022) [Dhawan2022] | LMC | TRGB | SNe Ia | ||
| 8 | Kenworthy (2022) [Kenworthy2022] | NGC 4258, LMC | Cepheids | — | ||
| 9 | Dhawan (2022) [Dhawan2023] | LMC | TRGB | SNe Ia | ||
| 10 | Dhawan (2022) [Dhawan2023] | NGC 4258, LMC, MW | Cepheids | SNe Ia | ||
| 11 | Scolnic (2023) [Scolnic2024] | LMC | TRGB | SNe Ia | ||
| 12 | de Jaeger (2023) [dejaeger2023] | LMC, MW | Cepheids | SNe II | ||
| 13 | Uddin (2023) [uddin2023] | Multiple | Combined | SNe Ia | ||
| 14 | Uddin (2023) [uddin2023] | Multiple | Combined | SNe Ia | ||
| 15 | Huang (2023) [Huang2024] | NGC 4258, LMC | Mira | SNe Ia | ||
| 16 | Li (2024) [Li2024b] | LMC, SMC | JAGB | SNe Ia | ||
| 17 | Chavez (2024) [chavez2025] | Cepheids | H II – | — | ||
| 18 | Boubel (2024) [Boubel2024] | Cepheids, TRGB | Tully–Fisher | — | ||
| 19 | Lee (2024) [Lee2024] | LMC | JAGB | SNe Ia | ||
| 20 | Freedman (2024) [Freedman2024] | LMC (DEB), MW | Cepheids | SNe Ia | ||
| 21 | Said (2024) [Said2025] | SBF | Fund. Plane | — | ||
| 22 | Zhang (2024)a [Zhang2024] | — | Pairwise Vel. | — | ||
| 23 | Jensen (2025) [Jensen2025] | LMC, MW | TRGB | — | ||
| 24 | Wojtak (2025) [Wojtak2025] | NGC 4258, LMC, MW | Cepheids | SNe Ia | ||
| 25 | Bhardwaj (2025) [Bhardwaj2025] | NGC 4258, LMC | Mira | SNe Ia | ||
| 26 | Newman (2025) [Newman2025] | LMC | TRGB | SNe Ia | ||
| 27 | Riess (2025)b [Riess2025] | NGC 4258, LMC, MW | Cepheids | SNe Ia | ||
| 28 | Kudritzki (2025) [kudritzki2025] | FGLR | Tully–Fisher | — | ||
| 29 | Wagner (2025)c [wagner2025] | — | TRGB | Infall |
@@
We adopt this measurement for all numerical values in this work. For comparisons, we use the SH0ES 2022 result [Riess2022] as reference baseline, given its established presence in the literature.
@@
A.2 One-Step Measurements: Extended Information
Table 10 provides extended information for all 54 One-Step measurements from Categories 2, 3, and 4. We include the physical observable, the cosmological probe, the model assumptions (which determine the category assignment), and the reported value with uncertainty.
@@ A.2 One-Step Measurements: Extended Information
Independent: Category 3 — Model-independent analysis (geometric, cosmographic, or direct )
@@
Strong Gravitational Lensing: 15 measurements (time delays from quasars, SNe, and clusters)
Gravitational Waves: 8 measurements (bright and dark sirens)
Fast Radio Bursts: 8 measurements (dispersion measure–redshift relation)
@@
Megamasers: 2 measurements (geometric disk distances)
CMB-based (no ): 6 measurements (SZ effect, scale)
Other: 11 measurements (EPM, kilonovae, gamma-ray attenuation, etc.)
@@
| Category 2: Local One-Step (CDM Assumption) — 33 measurements | |||||||
| 1 | 2 | Wu (2021) [Wu2022] | Dispersion measure, | FRBs | CDM | 0.0039–0.695 | |
| 2 | 2 | Abbott (2021) [Abbott2021] | GW strain amplitude | CBC dark sirens | CDM | ||
| 3 | 2 | Philcox (2022) [Philcox2022] | BAO scale, shape | Matter–rad. eq. | CDM | 0.38–0.61 | |
| 4 | 2 | Shajib (2023) [Shajib2023] | Time delays, dynamics | Lensed quasars | CDM | 0.657 | |
| 5 | 2 | Napier (2023) [Napier2023] | Time delays | Cluster lenses | CDM | 1.734–2.805 | |
| 6 | 2 | Kelly (2023) [Kelly2023] | Time delays | SN Refsdal | CDM | 1.49 | |
| 7 | 2 | Dominguez (2023) [Dominguez2023] | -ray opacity | Blazar spectra | CDM | ||
| 8 | 2 | Gao (2023) [gao2023] | DM– + SNe Ia | FRBs | CDM | 0.039–2.3 | |
| 9 | 2 | Liu (2023) [liu2023h] | Time delays | Cluster lensed QSO | CDM | ||
| 10 | 2 | Fung (2023) [Fung2023] | GW strain only | NSBH population | CDM | ||
| 11 | 2 | Martinez (2023) [Martinez2023] | Time delays | Lensed quasars | CDM | 1.734 | |
| 12 | 2 | Ballard (2023) [Ballard2023] | GW strain + galaxy | Dark siren | CDM | ||
| 13 | 2 | Moresco (2023) [Moresco2023] | Spectroscopic ages | Passive galaxies | CDM | 0.07–1.965 | |
| 14 | 2 | Alfradique (2024) [Alfradique2023] | GW strain + galaxy | Dark sirens | CDM | ||
| 15 | 2 | Grillo (2024) [Grillo2024] | Time delays | HFF clusters | CDM | 1.240–3.703 | |
| 16 | 2 | Pascale (2024) [Pascale2024] | Time delays | H0pe SN (JWST) | CDM | 1.78 | |
| 17 | 2 | Hernandez (2024) [Hernandez2024] | GW + BGP | Spectral sirens | CDM | ||
| 18 | 2 | Bom (2024) [Bom2024] | GW + kilonova | O4a events | CDM | 0.1–0.6 | |
| 19 | 2 | TDCOSMO XVI (2024) [TDCOSMO2024] | Time delays | WGD2038 | CDM | 0.777 | |
| 20 | 2 | Gao (2024) [Gao2025a] | DM– relation | Localized FRBs | CDM | 0.0008–1.016 | |
| 21 | 2 | Li (2024) [Li2024c] | Time delays (corrected) | Lensed quasars | CDM | ||
| 22 | 2 | Yang (2024) [Yang2024] | DM + scattering | FRBs | CDM | 0.034–1.016 | |
| 23 | 2 | Piratova-Moreno (2025) [PiratovMoreno2025] | DM– relation | FRBs (MLE) | CDM | 0.0008–1.354 | |
| 24 | 2 | Barua (2025) [Barua2025] | Maser dynamics | MCP sample | CDM | 0.002–0.034 | |
| 25 | 2 | Zhang (2025) [Zhang2025] | PRS association | FRBs | CDM | 0.03–0.25 | |
| 26 | 2 | Gao (2025) [Gao2025] | RM–PRS correlation | FRBs | CDM | 0.098–0.241 | |
| 27 | 2 | Loubser (2025) [Loubser2025] | Spectroscopic ages | BCGs | CDM | 0.3–0.7 | |
| 28 | 2 | Beirnaert (2025) [Beirnaert2025] | GW strain | Dark sirens (avg) | CDM | ||
| 29 | 2 | Bahr-Kalus (2025) [Bahr-Kalus2025] | Turnover scale + BAO | DESI | CDM | 0.4–3.1 | |
| 30 | 2 | Birrer (2025) [Birrer2025] | Time delays + kinematics | Multiple lenses | CDM | 0.082–0.745 | |
| 31 | 2 | Liu (2025) [liu2025a] | Strong grav. lensing | Multiple lenses | CDM | 1.49 | |
| 32 | 2 | Pierel (2025) [Pierel2025] | Time delays | SN Encore | CDM | 1.95 | |
| 33 | 2 | Paic (2025) [paic2025tdcosmoxxivmeasurementhubble] | Strong Lensing | SMARTS | CDM | 2.32 | |
| Category 3: Pure Local One-Step (Model Independent) — 15 measurements | |||||||
| 34 | 3 | Kuo (2013) [Kuo2013] | Maser positions, velocities | Megamaser disk | Independent | 0.034 | |
| 35 | 3 | Gao (2015) [Gao2015] | Maser positions, velocities | Megamaser disk | Independent | 0.028 | |
| 36 | 3 | Bulla (2022) [bulla2022] | GW + kilonova light curve | BNS merger | Independent | 0.01 | |
| 37 | 3 | Zhang (2022) [Zhang2022] | Spectroscopic ages | Passive galaxies | Independent | 0.009–2.545 | |
| 38 | 3 | Du (2023) [Du2023] | Time delays + GRBs | Lensed quasars | Independent | 0.654–2.375 | |
| 39 | 3 | Palmese (2023) [Palmese2023] | GW + afterglow | GW170817 | Independent | 0.01 | |
| 40 | 3 | Sneppen (2023) [Sneppen2023] | Photospheric velocity | Kilonova | Independent | 0.0986 | |
| 41 | 3 | Liu (2023) [liu2023a] | Lensing + SNe Ia | Combined | Independent | 0.01–2.3 | |
| 42 | 3 | Gonzalez (2024) [Gonzalez2024] | Gas fraction + SNe Ia | Clusters | Independent | 0.01–2.3 | |
| 43 | 3 | Li (2024) [Li2024a] | Time delays | SN Refsdal | Independent | 0.001–2.261 | |
| 44 | 3 | Jaiswal (2024) [Jaiswal2024] | RM lag | NGC 5548 | Independent | 0.017 | |
| 45 | 3 | Vogl (2024) [Vogl2024] | Photospheric expansion | SNe II | Independent | 0.01–0.04 | |
| 46 | 3 | Song (2025) [Song2025] | GW + lensing | Combined | Independent | ||
| 47 | 3 | Colaco (2025) [Colaco2025] | Lensing + SNe Ia | Joint | Independent | 0.001–2.261 | |
| 48 | 3 | Favale (2025) [Favale2025] | Ages + SNe Ia | CC + Pantheon+ | Independent | 0.07–1.965 | |
| Category 4: CMB Sound Horizon Free — 6 measurements | |||||||
| 49 | 4 | Reese (2003) [Reese2003] | SZ decrement, X-ray | Galaxy clusters | CMB–NSH | 0.02–0.83 | |
| 50 | 4 | Colaco (2023) [Colaco2023] | SZ + X-ray + SNe Ia | Clusters | CMB–NSH | 0.01–2.3 | |
| 51 | 4 | Pogosian (2024) [Pogosian2024] | BAO peak + BBN | DESI BAO | CMB–NSH | 0.1–4.16 | |
| 52 | 4 | Pogosian (2024) [Pogosian2024] | BAO peak + BBN | Pre-DESI BAO | CMB–NSH | 0.1–4.16 | |
| 53 | 4 | Bahr-Kalus (2025) [Bahr-Kalus2025] | Turnover scale + SNe | DESI + Pantheon+ | CMB–NSH | 0.4–3.1 | |
| 54 | 4 | Krolewski (2025) [Krolewski2025] | DESI DR1 + CMB | No sound horizon | CMB–NSH | 0.1–2.33 | |

### v2

@@ Acknowledgements.
This research was supported by COST Action CA21136 – Addressing observational tensions in cosmology with systematics and fundamental physics (CosmoVerse), supported by COST (European Cooperation in Science and Technology). We thank Adam Riess and Dan Scolnic for their detailed and constructive comments on measurement categorization, which helped us strengthen and clarify our analysis. We are grateful to Levon Pogosian for providing the unpublished DESI DR2 Sound-Horizon-Free result and to Helena García Escudero and Seyed Hamidreza Mirpoorian for sharing their work prior to publication. We thank Pilar Ruiz-Lapuente for useful correspondence regarding the JAGB calibration debate and for providing her “SNe Ia twins” measurement prior to its inclusion in our compilation. We also thank Yi-Ying Wang for pointing out relevant FRB and gravitational wave measurements, and Yi-min Zhang for sharing results on interacting dark energy models.Appendix A Complete Data Tables
@@
| 1. Distance Ladder | 30 | 72.73 | 0.39 | 0.72 | 67.8–77.0 |
| 2. Local (CDM) | 33 | 67.61 | 0.96 | 0.80 | 54.6–86.2 |
| 3. Pure Local | 16 | 71.03 | 0.69 | 0.77 | 65.9–75.5 |
| 4. CMB–SHF | 9 | 69.07 | 0.44 | 0.89 | 61–70.8 |
| Combined DLI–SHF (2+3+4) | 58 | 69.37 | 0.34 | 0.95 | 54.6–86.2 |
@@ A.1 Distance Ladder Measurements: Extended Information
Table 10 provides extended information for all 30 Distance-Ladder-dependent measurements. We include the geometric anchor(s) employed, the primary distance indicator, the secondary distance indicator (if applicable), and the reported value with uncertainty.
@@
| 1 | Huang (2019) [78] | NGC 4258, MW, LMC | Mira | SNe Ia | ||
| 2 | de Jaeger (2020) [42] | LMC, MW | Cepheids/TRGB | SNe II | ||
| 3 | Kourkchi (2020) [92] | Cepheids, TRGB | Tully–Fisher | — | ||
| 4 | Khetan (2021) [90] | Cepheids, TRGB | SBF | SNe Ia | ||
| 5 | Blakeslee (2021) [22] | NGC 4258, LMC | SBF | SNe Ia | ||
| 6 | Freedman (2021) [63] | LMC (DEB), MW | TRGB | SNe Ia | ||
| 7 | Dhawan (2022) [47] | LMC | TRGB | SNe Ia | ||
| 8 | Kenworthy (2022) [88] | NGC 4258, LMC | Cepheids | — | ||
| 9 | Dhawan (2022) [48] | LMC | TRGB | SNe Ia | ||
| 10 | Dhawan (2022) [48] | NGC 4258, LMC, MW | Cepheids | SNe Ia | ||
| 11 | Scolnic (2023) [161] | LMC | TRGB | SNe Ia | ||
| 12 | de Jaeger (2023) [41] | LMC, MW | Cepheids | SNe II | ||
| 13 | Uddin (2023) [180] | Multiple | Combined | SNe Ia | ||
| 14 | Uddin (2023) [180] | Multiple | Combined | SNe Ia | ||
| 15 | Huang (2023) [79] | NGC 4258, LMC | Mira | SNe Ia | ||
| 16 | Li (2024) [98] | LMC, SMC | JAGB | SNe Ia | ||
| 17 | Chavez (2024) [34] | Cepheids | H II – | — | ||
| 18 | Boubel (2024) [24] | Cepheids, TRGB | Tully–Fisher | — | ||
| 19 | Lee (2024) [96] | LMC | JAGB | SNe Ia | ||
| 20 | Freedman (2024) [62] | LMC (DEB), MW | Cepheids | SNe Ia | ||
| 21 | Said (2024) [156] | SBF | Fund. Plane | — | ||
| 22 | Zhang (2024)a [193] | — | Pairwise Vel. | — | ||
| 23 | Jensen (2025) [83] | LMC, MW | TRGB | — | ||
| 24 | Wojtak (2025) [186] | NGC 4258, LMC, MW | Cepheids | SNe Ia | ||
| 25 | Bhardwaj (2025) [20] | NGC 4258, LMC | Mira | SNe Ia | ||
| 26 | Newman (2025) [117] | LMC | TRGB | SNe Ia | ||
| 27 | Riess (2025)b [149] | NGC 4258, LMC, MW | Cepheids | SNe Ia | ||
| 28 | Kudritzki (2025) [94] | FGLR | Tully–Fisher | — | ||
| 29 | Wagner (2025)c [183] | — | TRGB | Infall | ||
| 30 | Lapuente (2025) [155] | — | M 101, NGC 5643, NGC 7250, M 66 | SNe Ia (twins) |
@@
We adopt this measurement for all numerical values in this work. For comparisons, we use the SH0ES 2022 result [150] as reference baseline, given its established presence in the literature.
@@
A.2 Distance Ladder Independent/Sound Horizon Free (DLI–SHF) Measurements: Extended Information
Table 11 provides extended information for all 58 DLI–SHF measurements from Categories 2, 3, and 4. We include the physical observable, the cosmological probe, the model assumptions (which determine the category assignment), and the reported value with uncertainty.
@@ A.2 Distance Ladder Independent/Sound Horizon Free (DLI–SHF) Measurements: Extended Information
Independent: Category 3 — Cosmological-model-independent analysis (geometric, cosmographic, or direct )
@@
Strong Gravitational Lensing: 18 measurements (time delays from quasars, SNe, and clusters)
Gravitational Waves: 9 measurements (bright and dark sirens)
Fast Radio Bursts: 7 measurements (dispersion measure–redshift relation)
@@
Megamasers: 3 measurements (geometric disk distances)
CMB-based (no ): 9 measurements (SZ effect, scale, etc.)
Other: 8 measurements (EPM, kilonovae, gamma-ray attenuation, etc.)
@@
| Category 2: Local (CDM Assumption) — 33 measurements | |||||||
| 1 | 2 | Wu (2021) [189] | Dispersion measure, | FRBs | CDM | 0.0039–0.695 | |
| 2 | 2 | Abbott (2021) [102] | GW strain amplitude | CBC dark sirens | CDM | ||
| 3 | 2 | Shajib (2023) [163] | Time delays, dynamics | Lensed quasars | CDM | 0.657 | |
| 4 | 2 | Napier (2023) [116] | Time delays | Cluster lenses | CDM | 1.734–2.805 | |
| 5 | 2 | Kelly (2023) [87] | Time delays | SN Refsdal | CDM | 1.49 | |
| 6 | 2 | Dominguez (2023) [53] | -ray opacity | Blazar spectra | CDM† | ||
| 7 | 2 | Gao (2023) [67] | DM– + SNe Ia | FRBs | CDM | 0.039–2.3 | |
| 8 | 2 | Liu (2023) [104] | Time delays | Cluster lensed QSO | CDM | ||
| 9 | 2 | Fung (2023) [64] | GW strain only | NSBH population | CDM | ||
| 10 | 2 | Martinez (2023) [113] | Time delays | Lensed quasars | CDM | 1.734 | |
| 11 | 2 | Ballard (2023) [14] | GW strain + galaxy | Dark siren | CDM | ||
| 12 | 2 | Moresco (2023) [115] | Spectroscopic ages | Passive galaxies | CDM | 0.07–1.965 | |
| 13 | 2 | Alfradique (2024) [11] | GW strain + galaxy | Dark sirens | CDM | ||
| 14 | 2 | Grillo (2024) [75] | Time delays | HFF clusters | CDM | 1.240–3.703 | |
| 15 | 2 | Pascale (2024) [125] | Time delays | H0pe SN (JWST) | CDM | 1.78 | |
| 16 | 2 | Hernandez (2024) [110] | GW + BGP | Spectral sirens | CDM | ||
| 17 | 2 | Bom (2024) [23] | GW + kilonova | O4a events | CDM | 0.1–0.6 | |
| 18 | 2 | TDCOSMO XVI (2024) [188] | Time delays | WGD2038 | CDM | 0.777 | |
| 19 | 2 | Gao (2024) [65] | DM– relation | Localized FRBs | CDM | 0.0008–1.016 | |
| 20 | 2 | Li (2024) [99] | Time delays (corrected) | Lensed quasars | CDM | ||
| 21 | 2 | Yang (2024) [190] | DM + scattering | FRBs | CDM | 0.034–1.016 | |
| 22 | 2 | Piratova-Moreno (2025) [140] | DM– relation | FRBs (MLE) | CDM | 0.0008–1.354 | |
| 23 | 2 | Barua (2025) [17] | Maser dynamics | MCP sample | CDM | 0.002–0.034 | |
| 24 | 2 | Zhang (2025) [195] | PRS association | FRBs | CDM | 0.03–0.25 | |
| 25 | 2 | Gao (2025) [69] | RM–PRS correlation | FRBs | CDM | 0.098–0.241 | |
| 26 | 2 | Loubser (2025) [107] | Spectroscopic ages | BCGs | CDM | 0.3–0.7 | |
| 27 | 2 | Beirnaert (2025) [18] | GW strain | Dark sirens (avg) | CDM | ||
| 28 | 2 | Bahr-Kalus (2025) [13] | Turnover scale + BAO | DESI | CDM | 0.4–3.1 | |
| 29 | 2 | Birrer (2025) [175] | Time delays + kinematics | Multiple lenses | CDM | 0.082–0.745 | |
| 30 | 2 | Liu (2025) [105] | Strong grav. lensing | Multiple lenses | CDM | 1.49 | |
| 31 | 2 | Pierel (2025) [138] | Time delays | SN Encore | CDM | 1.95 | |
| 32 | 2 | Paic (2025) [121] | Strong Lensing | SMARTS | CDM | 2.32 | |
| 33 | 2 | Oliveira (2026) [12] | 3×2pt + Standard Sirens | Shear/positions + GW strain | CDM | 0.1-2 | |
| Category 3: Pure Local (Cosmological-Model Independent) — 16 measurements | |||||||
| 34 | 3 | Kuo (2013) [95] | Maser positions, velocities | Megamaser disk | Independent | 0.034 | |
| 35 | 3 | Gao (2015) [66] | Maser positions, velocities | Megamaser disk | Independent | 0.028 | |
| 36 | 3 | Bulla (2022) [26] | GW + kilonova light curve | BNS merger | Independent | 0.01 | |
| 37 | 3 | Zhang (2022) [192] | Spectroscopic ages | Passive galaxies | Independent | 0.009–2.545 | |
| 38 | 3 | Du (2023) [55] | Time delays + GRBs | Lensed quasars | Independent | 0.654–2.375 | |
| 39 | 3 | Palmese (2023) [122] | GW + afterglow | GW170817 | Independent | 0.01 | |
| 40 | 3 | Sneppen (2023) [171] | Photospheric velocity | Kilonova | Independent | 0.0986 | |
| 41 | 3 | Liu (2023) [103] | Lensing + SNe Ia | Combined | Independent | 0.01–2.3 | |
| 42 | 3 | Gonzalez (2024) [74] | Gas fraction + SNe Ia | Clusters | Independent | 0.01–2.3 | |
| 43 | 3 | Li (2024) [101] | Time delays | SN Refsdal | Independent | 0.001–2.261 | |
| 44 | 3 | Jaiswal (2024) [81] | RM lag | NGC 5548 | Independent | 0.017 | |
| 45 | 3 | Vogl (2024) [182] | Photospheric expansion | SNe II | Independent | 0.01–0.04 | |
| 46 | 3 | Song (2025) [172] | GW + lensing | Combined | Independent | ||
| 47 | 3 | Colaco (2025) [37] | Lensing + SNe Ia | Joint | Independent | 0.001–2.261 | |
| 48 | 3 | Du (2025) [54] | CC+BAO+ SNe Ia | IDL+ MAPAge | Independent | 0.001–2.5 | |
| 49 | 3 | Favale (2025) [60] | Ages + SNe Ia | CC + Pantheon+ | Independent | 0.07–1.965 | |
| Category 4: CMB Sound Horizon Free (CMB–SHF) — 9 measurements | |||||||
| 50 | 4 | Reese (2003) [147] | SZ decrement, X-ray | Galaxy clusters | CMB–NSH | 0.02–0.83 | |
| 51 | 4 | Philcox (2022) [137] | BAO scale, shape | Matter–rad. eq. | CMB–NSH | 0.2–0.75 | |
| 52 | 4 | Colaco (2023) [36] | SZ + X-ray + SNe Ia | Clusters | CMB–NSH | 0.01–2.3 | |
| 53 | 4 | Pogosian (2024) [144] | BAO peak + BBN | Pre-DESI BAO | CMB–NSH | 0.1–4.16 | |
| 54 | 4 | Pogosian (2025) [145] | BAO peak + BBN | DESI DR BAO | CMB–NSH | 0.1–4.16 | |
| 55 | 4 | Bahr-Kalus (2025) [13] | Turnover scale + SNe | DESI + Pantheon+ | CMB–NSH | 0.4–3.1 | |
| 56 | 4 | Garcia (2025) [70] | BAO+ +CMB lensing | DESI+APS+ DES 3×2pt | CMB–NSH | 0.1-4.2 | |
| 57 | 4 | Zaborowski [191] | BAO+ | Clust.+ BAO+ Lym.- | CMB–NSH | 0.1-4.2 | |
| 58 | 4 | Krolewski (2025) [93] | DESI DR1 + CMB | Energy densities | CMB–NSH | 0.1–2.33 | |